### v1

@@ 2.1.1 Kitagawa (2015), and Mourifié and Wan (2017) testable implications
One significant difference between the testable implication of Kitagawa (2015) and Mourifié and Wan (2017) and ours is we do not assume a prior direction. To illustrate this point, suppopse $\mathcal{Z}=\{z_{1},...,z_{K}\}$ and suppose we assume one of the $K!$ potential directions as:
$D_{z_{K}}\geq D_{z_{K-1}}\geq...\geq D_{z_{1}}$
meaning that type $z_{K}$ judge is less lenient than type $z_{K-1}$ judge, which, in turn, is less lenient than $z_{K-2}$, $z_{K-3}$, $\cdots,z_{1}$ judge. Assumptions 2.1, 2.2 and 2.3, plus the above-imposed direction, imply the following testable implications studied in Sun (2023):
@@
[…]
A key point to note is the above implications restrict $F_{Y,D|Z}(y,d|z)$ while the testable implications in Theorem 2.1(i) instead restrict $F_{Y,D|P}(y,d|p)$. In the first case, the induced direction of inequalities is with respect to the observed judge type $Z$, while in our case, the inequalities are with respect to the propensity score $P$, which is obtained without imposing a prior direction. Also, noteworthy is if one takes $y=-\infty$ and $y^{\prime}=\infty$, the testable implications in Theorem 2.1(i) no longer have any empirical content. But, the testable implications with an ex-ante monotonicity direction still restrict the propensity scores and the judges’ types, i.e., $P$, and $Z$, such that
@@ 2.1.2 Frandsen, Lefgren, and Leslie (2023)’s test
FLL proposes a set of testable implications for Assumptions 2.1, 2.2 and 2.3. Their testable implication has sound features of not relying on the ex-ante specified direction of monotonicity and assessing all the assumptions jointly. Their testable implication is not sharp and can fail to screen some observable violations of the judge leniency design. To see this, consider any integrable function $g(\cdot):\mathcal{Y}\rightarrow\mathbb{R}$, and let $p\neq p^{\prime}\in\mathcal{P}$. Under Assumptions 2.1, 2.2 and 2.3, we can derive the following equality:
@@
One can easily observe that the testable restrictions in (2.4) and (2.5) could be violated, whereas the restriction used by FLL, i.e. inequality (2.3) still holds. These discordant implications confirm the concern about developing a statistical test based on non-sharp restrictions. Hence, implementing FLL’s statistical testing procedure based on inequalities (2.4) or (2.5) could provide a different result compared to their test based on inequality (2.3) alone.
Another evident reason why FLL’s testing approach cannot exhaust all violations of the judge leniency design is they only focus on $g(Y)=Y$ whereas the inequality in (2.3) should hold for any integrable function $g$ and for any pair $p\neq p^{\prime}\in\mathcal{P}$. $g(Y)=Y$ is not a sufficient class of functions to screen all violations of the model.
Finally, we note our testable implications in Theorem 2.1 do not rely on the known support of $g(Y)$, whereas to test inequality (2.3), one needs to know the bounds of the support $(U_{g},L_{g})$. If the support of $g(Y)$ is unbounded, i.e. $U_{g}=+\infty$ and $L_{g}=-\infty$, then the testable implication in (2.3) holds trivially and does not have any power in detecting violations to the identification assumptions.
@@ 3 Testing Procedures
We first present our baseline semi-nonparametric test in Section 3.1 without the presence of control variables $X$. For this test, we do not make any functional form or distributional assumptions on potential outcomes. For the propensity score, we follow the common practice in the literature to employ a parametric model so that $P(z)=P(z,\theta_{0})$ for all $z\in\mathcal{Z}$ and for a finite-dimensional parameter vector $\theta_{0}\in\Theta$. Popular choices include the Probit or Logit model with a linear index $z^{\prime}\theta_{0}$ (see, for instance, Carneiro, Heckman, and Vytlacil, 2011; Kowalski, 2016, among many others). We also want to emphasize when the instrument variable $Z$ is the judge’s identity, we do not need to impose any parametric assumptions on the propensity score. It can be estimated by the sample average of $D$ conditioning on each judge. In this case, our test is indeed nonparametric.
@@ 3.1 A Semi-nonparametric test
For the convenience of the exposition, we re-state the testable implications as the null hypothesis $H_{0}$. That is, for all $p_{1}\geq p_{2}\text{ with }p_{1},p_{2}\in\mathcal{P}\text{ and all }y,y^{\prime}\in\mathcal{Y}$,
@@
The alternative hypothesis $H_{1}$ is then inequality (3.1) or (3.2) fails to hold for some $(p_{1},p_{2})$ and $(y,y^{\prime})$. Without loss of generality, we assume the support of $Y$ is $[0,1]$. [^8] We can always apply a transformation to ensure the support of $Y$ is $[0,1]$. If $Y$ has a finite support $[a,b]$, we can apply an affine transformation $\tilde{Y}=(Y-a)/(b-a)$. If $Y$’s support is the whole real line, we can apply standard normal CDF after rescaling and recentering: $\tilde{Y}=\Phi\left(\frac{Y-\bar{Y}}{\hat{std}(Y)}\right)$, where $\bar{Y}$ is the sample average and $\hat{std}(Y)$ is the sample standard deviation. Testing inequalities (3.1) and (3.2) involves two features; first, it is a set of inequality restrictions defined on conditional moments where the conditioning variable is possibly continuous. We deal with the first difficulty by employing the method of Hsu, Liu, and Shi (2019) to transform them into an equivalent set of restrictions on unconditional moments. The second feature is the conditioning variable $P$ is not directly observed from the data. For this latter feature, we match our testing procedure with most of the empirical practice and estimate the propensity score parametrically in the first step. We derive the new influence functions and show the first-stage estimation error is properly accounted for.
@@
${\mathcal{L}}_{Y}=\left\{(y,r_{y}):~{}r_{y}=q_{y}^{-1},~{}\text{ }q_{y}\cdot r_{y}\in\{0,1,2,\cdots,(q_{y}-1)\}\text{ for }q_{y}=1,2,\cdots,\right\}.$
@@
In the expression of $\nu_{d}(\ell)$, the propensity score $P(Z,\theta_{0})$ is unknown, but can be replaced by its estimate $\hat{P}\equiv P(Z,\hat{\theta})$, where $\hat{\theta}$ is the MLE. Under this parameterization, we also write $\nu_{d}(\ell)$ as $\nu_{d}(\ell,\theta_{0})$ when it causes no confusion. 3.1 below summarizes the semi-nonparametric test’s implementation procedure. Please see Appendix A for detailed equations and expressions.
@@
    Specify large integers $Q_{Y}$ and $Q_{P}$, and create a coarser version $\mathcal{L}_{Q}$ of $\mathcal{L}$ set by limiting $q_{y}=1,2,\cdots,Q_{Y}$ and $q_{p}=1,2,\cdots,Q_{P}$.
@@
    Estimate the propensity score by $\hat{P}_{i}=P(Z_{i},\hat{\theta})$, where $P(z,\theta)$ parameterized up to $\theta$ and $\hat{\theta}$ is the maximum likelihood estimator.
@@
    For each $\ell\in\mathcal{L}_{Q}$, construct estimates $\hat{\nu}_{1}(\ell)$ and $\hat{\nu}_{0}(\ell)$ using Equations A.6 and A.7.
@@
    Choose a constant $\epsilon>0$. For each $\ell\in\mathcal{L}_{Q}$, construct $\hat{\sigma}^{2}_{d,\epsilon}(\ell)=\max\{\hat{\sigma}^{2}_{d}(\ell),\epsilon\}$, where $\hat{\sigma}^{2}_{d}(\ell)$ is a consistent estimator for the asymptotic variance of $\sqrt{n}(\widehat{\nu}_{d}(\ell,\widehat{\theta})-\nu_{d}(\ell,\theta_{0}))$ and is defined below in Equation 3.8.
  5. 5.
  6. 6.
  7. 7.
@@
      Estimate propensity score for each bootstrap iteration $\hat{P}_{i}^{b}=P(Z_{i},\hat{\theta}^{b})$, where $\hat{\theta}^{b}$ is the maximum likelihood estimator.
@@
      Obtain $\hat{\nu}_{d}^{b}(\ell)$, $d=0,1$, for each bootstrap iteration using Equations A.9 and A.8.
    4. (d)
      Calculate the quantity
@@
Theorem 3.2 shows that the test $\phi_{n}$ has its size controlled asymptotically and is consistent. The proof for Theorem 3.2 is collected in Section C of the online supplementary material. We also list all the technical conditions in that section for exposition purposes.
@@
Suppose Assumptions C.2, C.3, C.4 and C.1 in Appendix D are satisfied. Let $\alpha\in(0,1/2)$ be the pre-chosen significance level.
@@
    $\limsup_{n\rightarrow\infty}\mathbb{P}(\phi_{n}=1|H_{0})=1.$ (3.10)
@@ 3.2 A semiparametric test with covariates dimension reduction
$Z\perp(Y_{0}(z),Y_{1}(z),D(z);z\in\mathcal{Z})|X=x$ for all $x\in\mathcal{X}$
@@
When Assumptions 3.1, 2.2 and 3.2 hold, the testable implications can be written as follows. For all $x\in\mathcal{X}$, $p_{1},p_{2}\in\mathcal{P}$ and $p_{1}\leq p_{2}$, and all $y,y^{\prime}\in\mathcal{Y}$
@@
If $X$ is discrete and $\mathcal{X}$ only contains a relatively small number of values, we can implement the procedure in Section 3.1 for each subsample defined by those support points, and control the familywise error rate by certain multiple testing procedure, e.g. Holm (1979). If $X$ contains a small number of continuous variables, we can also follow the same procedure as in Section 3.1 but adding cubes for $X$. For example, it implies $\nu_{1}(\ell,x)\leq 0$ for all $(\ell,x)$ where
@@
$\nu_{0}(\ell)\equiv\mathbb{E}[(D-1)1\{y\leq Y\leq y+r_{y}\}1\{x\leq X\leq x+r_{x}\}1\{p_{2}\leq P\leq p_{2}+r_{p}\}]\\ \times\mathbb{E}[1\{x\leq X\leq x+r_{x}\}1\{p_{1}\leq P\leq p_{1}+r_{p}\}]-\mathbb{E}[1\{x\leq X\leq x+r_{x}\}1\{p_{2}\leq P\leq p_{2}+r_{p}\}]\\ \times{E}[(D-1)1\{y\leq Y\leq y+r_{y}\}1\{x\leq X\leq x+r_{x}\}1\{p_{1}\leq P\leq p_{1}+r_{p}\}],$
and $r_{x}$ is similarly defined as $r_{p}$ and $r_{y}$. The implementation follows analously from 3.1.
@@
The advantage of using (3.13) and (3.14) is both inequalities are only conditional on the scalar-valued propensity score. The effect of covariates has been filtered out by constructing a new outcome variable $\tilde{Y}$. Under the null hypothesis of the model being correctly specified, parameters $\beta_{0}$ and $\beta_{1}$ can be estimated by partial linear regression of $Y$ on $X$ and propensity score $P$ separately for the sample of $D=1$ and $D=0$; see for instance Carneiro and Lee (2009); Carneiro, Heckman, and Vytlacil (2010); Kowalski (2016). Specifically, for the potential outcome $Y_{1}$, we have
$\mathbb{E}[Y_{1}|X=x,V=p]=x^{\prime}\beta_{1}+\underbrace{\mathbb{E}[\alpha_{1}+U_{1}|X=x,V=p]}_{k_{1}(p)},$
and $\beta_{1}$ and $k_{1}(\cdot)$ can be estimated by the following partial linear regression:
$\mathbb{E}[Y|X=x,P=p,D=1]=x^{\prime}\beta_{1}+K_{1}(p),\quad k_{1}(p)=K_{1}(p)+pK_{1}^{\prime}(p).$
Likewise, for potential outcome $Y_{0}$, we have
$\mathbb{E}[Y_{0}|X=x,V=p]=x^{\prime}\beta_{0}+\underbrace{\mathbb{E}[\alpha_{0}+U_{0}|X=x,V=p]}_{k_{0}(p)}$
and $\beta_{0}$ and $k_{0}(\cdot)$ can be estimated by the following partial linear regression:
$\mathbb{E}[Y|X=x,P=p,D=0]=x^{\prime}\beta_{0}+K_{0}(p),\quad k_{0}(p)=K_{0}(p)-(1-p)K_{0}^{\prime}(p)$
The following algorithm summarizes the steps for implementation.
@@
    The procedure starts with estimated propensity score $\hat{P}_{i}=P(Z_{i},X_{i},\hat{\theta})$ using Equation E.1.
@@
    Once $\hat{\beta}_{1}$ and $\hat{\beta}_{0}$ are obtained, one can construct estiamtes for ${\widetilde{Y}}_{i}=Y_{i}-X_{i}^{\prime}(D_{i}\hat{\beta}_{1}+(1-D_{i})\hat{\beta}_{0})$

### v2

@@ 2.1.1 Kitagawa (2015), and Mourifié and Wan (2017) testable implications
One significant difference between the testable implication of Kitagawa (2015) and Mourifié and Wan (2017) and ours is we do not assume a prior direction. To illustrate this point, suppose $\mathcal{Z}=\{z_{1},...,z_{K}\}$ and suppose we assume one of the $K!$ potential directions as: D_z_K≥D_z_K-1 ≥…≥D_z_1 meaning that type $z_{K}$ judge is less lenient than type $z_{K-1}$ judge, which, in turn, is less lenient than $z_{K-2}$, $z_{K-3}$, $\cdots,z_{1}$ judge. Given this imposed ordering, Assumptions 2.1, 2.2 and 2.3 imply the following testable implications studied in Sun (2023):
@@
 $~~~\text{ for all }k\in\{1,...,K-1\}\text{ and }y,y^{\prime}\in\mathcal{Y}.$
A key point to note is that the above implications restrict $F_{Y,D|Z}(y,d|z)$ while the testable implications in Theorem 2.1(i) instead restrict $F_{Y,D|P}(y,d|p)$. In the first case, the induced direction of inequalities is with respect to the observed judge type $Z$, while in our case, the inequalities are with respect to the propensity score $P$, which is obtained without imposing a prior direction. Also, noteworthy is that if one takes $y=$-$\infty$ and $y^{\prime}=\infty$, the testable implications in Theorem 2.1(i) no longer have any empirical content. But, the testable implications with an ex-ante monotonicity direction still restrict the propensity scores and the judges’ types, i.e., $P$, and $Z$, such that
@@ 2.1.2 Frandsen, Lefgren, and Leslie (2023)’s test
FLL proposes a set of testable implications for Assumptions 2.1, 2.2 and 2.3. Their testable implication has sound features of not relying on the ex-ante specified direction of monotonicity and assessing all the assumptions jointly. Their testable implication, however, is not sharp and can fail to screen some non-negligible observable violations of the judge leniency design. To see this, consider any integrable function $g(\cdot):\mathcal{Y}\rightarrow\mathbb{R}$, and let $p\neq p^{\prime}\in\mathcal{P}$. Under Assumptions 2.1, 2.2 and 2.3, we can derive the following equality:
@@
One can easily observe that the testable restrictions in (2.3) and (2.4) could be violated, whereas the restriction used by FLL, i.e. inequality (2.2) still holds. Hence, implementing FLL’s statistical testing procedure based on inequalities (2.3) or (2.4) could provide a different result compared to their test based on inequality (2.2) alone. These discordant implications confirm the concern about developing a statistical test based on non-sharp restrictions. Example 2.1 provides a concrete numerical example.
Example 2.1
Consider the potential outcome model:
$\begin{cases}Y=Y_{1}D+Y_{0}(1-D),\\ D=1\left\{P\geq V\right\}.\end{cases}$
Suppose $V$ is independent of $(Y_{1},Y_{0},P)$. However, $(Y_{1},Y_{0})$ and $P$ are dependent:
[…]
Therefore, the randomization assumption is violated, but the monotonicity and exclusion conditions are met. In this case, $Y_{d}$ is binary and $U_{g}=1$ and $L_{g}=0$, so we also take $g(\cdot)$ to be the identity function without loss of generality.
For this DGP, we can show that for any $p^{\prime}\in(0,1)$ and $p\in(0,1)$, $W(g(Y),p,p^{\prime})=1$. Hence, FLL’s testable implication (inequality 2.2 above) always holds and has no power to detect the violation. There is missing information. For example, when $p^{\prime}>p>\frac{1}{2}$, we can verify that $W(g(YD),p,p^{\prime})=p^{\prime}+p>1\equiv U_{g}$. Therefore, condition (2.3) is violated. Please see derivation details in Section B.4.
On the other hand, our testable implication can capture such a violation. To see this, note
$\mathbb{E}[YD|P=p]=\mathbb{E}[Y_{1}|P=p]p=\begin{cases}p\;\;\text{ if }p<\frac{1}{2},\\ p^{2}\;\;\text{ if }p\geq\frac{1}{2}.\\ \end{cases}.$
[…]
The intuition behind Example 2.1 is not pathological and is reflected in the derivation in Section B.4. Because $\mathbb{E}[Y|P=p]=\mathbb{E}[Y_{1}D|P=p]+\mathbb{E}[Y_{0}(1-D)|P=p]$, it is possible the violations on the $\mathbb{E}[Y_{1}D|P=p]$ and $\mathbb{E}[Y_{0}(1-D)|P=p]$ “cancel” out. As a consequence, the quantity $\mathbb{E}[Y|P=p]$ provides no power to detect violations in these cases.
Another evident reason why FLL’s implications cannot exhaust all violations of the judge leniency design is that they only focus on $g(Y)=Y$, whereas the inequality in (2.2) should hold for any integrable function $g$ and for any pair $p\neq p^{\prime}\in\mathcal{P}$. $g(Y)=Y$ is not a sufficient class of functions to screen all violations of the model.
Finally, we note our testable implications in Theorem 2.1 do not rely on the known support of $g(Y)$, whereas to test inequality (2.2), one needs to know the bounds of the support $(U_{g},L_{g})$. If the support of $g(Y)$ is unbounded, i.e., $U_{g}=+\infty$ and $L_{g}=$-$\infty$, then the testable implication in (2.2) holds trivially and FLL’s test does not have any power in detecting violations to the identification assumptions.
@@ 3 Testing Procedures
In this section, we construct tests based on Theorem 2.1. For the defendant $i\in\{1,2,\cdots,n$}, researchers observed a vector $(Y_{i},D_{i},Z_{i},X_{i})$, where $Y_{i}$, $D_{i}$, $Z_{i}$, and $X_{i}$ represent his/her observed outcome, observed treatment status, the vector of characteristics of the judge that $i$ was assigned to, and the vector of additional control variables, respectively.
We first present our baseline semi-nonparametric test in Section 3.1 without the presence of control variables $X_{i}$. For this test, we make no functional form or distributional assumptions about potential outcomes. We do need to estimate the propensity score $P(z)\equiv\mathbb{P}(D_{i}=1|Z_{i}=z)$ first, for which our procedure can accommodate different data scenarios. If $Z_{i}$ contains continuous variables, we follow the common practice in the literature to employ a parametric model so that $P(z)=P(z,\theta_{0})$ for all $z\in\mathcal{Z}$ and for a finite-dimensional parameter vector $\theta_{0}\in\Theta$. Popular choices include the Probit or Logit model with a linear index $z^{\prime}\theta_{0}$ (see, for instance, Carneiro, Heckman, and Vytlacil, 2011; Kowalski, 2016, among many others). [^9] When $Z$ is continuous, the rejection result of our semi-nonparametric test can be interpreted as rejecting the joint assumption of the judge leniency design and the parametric form imposed on the propensity score. In our simulation studies, we always keep the propensity score correctly specified. In these studies, therefore, the rejection shows the power of our test to reject false judge leniency assumptions. When $Z_{i}$ only contains discrete variables, such as judge’s gender, we can estimate $\mathbb{P}(D_{i}=1|Z_{i}=z)$ by the sample averages of $D$ conditioning on each possible value of $z\in\mathcal{Z}$. In this case, our test is indeed nonparametric.
We should emphasize that, in both cases above, we do not require any knowledge of the identity of judges, nor do we need the number of defendants handled by each judge to diverge to infinity. For example, when $Z$ is gender, we only need the number of defendants for each judge’s gender to go to infinity. This can happen when the number of judges is large, but each judge handles a finite number (or even one) of defendants. Therefore, our test can also be applied to other empirical contexts than the judge leniency design. There is another scenario in which $Z_{i}$ is the judge’s identity. Suppose the number of defendants handled by each judge is large, as in our empirical application. In this case, we can also consistently estimate judge $j$’s propensity score $\mathbb{P}(D=1|Z_{i}=j)$ by the sample frequency estimator for each judge $j$, regardless of whether the number of judges is small or large.
In practice, the number of defendants that a judge handles can be small. In this case, one can not estimate $\mathbb{P}(D=1|Z_{i}=j)$ consistently without additional assumptions and conventional inference methods can be invalid. This phenomenon has received attention from the literature, see discussions in Jochmans (2023), Ren (2024), Sithole (2024), and Yap (2024). These papers, however, focus on inference on the parameters instead of testing model specification. To account for data scenarios with small numbers of defendants per judge, we design a test that does not require a consistent estimator for the propensity score and controls the size at any sample sizes for the case of a binary outcome. To the best of our knowledge, this test and FLL’s finite sample test are the only ones for testing judge leniency design specification with finite samples, and both focus on binary outcome variables. Our test uses upper bounds of the null distribution to calculate the critical values, and hence is very easy to implement. It only requires simulating Bernoulli random variables, and no nonlinear optimization is involved. For the purpose of exposition, we collect the finite sample test in Appendix C and focus on the cases in which the propensity score can be consistently estimated in this section.
@@ 3.1 A Semi-nonparametric test
For the convenience of the exposition, we restate the testable implications as the null hypothesis $H_{0}$. That is, for all $p_{1}\geq p_{2}\text{ with }p_{1},p_{2}\in\mathcal{P}\text{ and all }y,y^{\prime}\in\mathcal{Y}$,
@@
The alternative hypothesis $H_{1}$ is then inequality (3.1) or (3.2) fails to hold for some $(p_{1},p_{2})$ and $(y,y^{\prime})$. Without loss of generality, we assume the support of $Y$ is [0,1][0,1]. [^10] We can always apply a transformation to ensure the support of $Y$ is $[0,1]$. If $Y$ has a finite support $[a,b]$, we can apply an affine transformation $\tilde{Y}=(Y-a)/(b-a)$. If $Y$’s support is the whole real line, we can apply standard normal CDF after rescaling and re-centering: $\tilde{Y}=\Phi\left(\frac{Y-\bar{Y}}{\hat{std}(Y)}\right)$, where $\bar{Y}$ is the sample average and $\hat{std}(Y)$ is the sample standard deviation. Testing inequalities (3.1) and (3.2) involves two features; first, it is a set of inequality restrictions defined on conditional moments where the conditioning variable is possibly continuous. We deal with the first difficulty by employing the method of Hsu, Liu, and Shi (2019) to transform them into an equivalent set of restrictions on unconditional moments. The second feature is that the conditioning variable $P$ is not directly observed from the data. We derive the new influence functions and show that the first-stage estimation error is properly accounted for.
@@
${\mathcal{L}}_{Y}=\left\{(y,r_{y}):~r_{y}=q_{y}^{-1},~\text{ }q_{y}\cdot y\in\{0,1,2,\cdots,(q_{y}-1)\}\text{ for }q_{y}=1,2,\cdots,\right\}.$
@@
[…]
@@
    Specify integers $Q_{Y}$ and $Q_{P}$, and create a coarser version $\mathcal{L}_{Q}$ of $\mathcal{L}$ set by limiting $q_{y}=1,2,\cdots,Q_{Y}$ and $q_{p}=1,2,\cdots,Q_{P}$.
@@
    Compute the estimator for the propensity score $\hat{P}_{i}$, as detailed in Appendix A.
@@
    For each $\ell\in\mathcal{L}_{Q}$, construct estimates $\hat{\nu}_{1}(\ell)$ and $\hat{\nu}_{0}(\ell)$ as sample analogs of Equations 3.3 and 3.4, as detailed in Equations A.5 and A.6.
@@
      Estimate propensity score for each bootstrap iteration $\hat{P}_{i}^{b}$, defined in Equation A.9.
@@
      Obtain $\hat{\nu}_{d}^{b}(\ell)$, $d=0,1$, for each bootstrap iteration using Equations A.11 and A.10.
  5. 5.
    Compute the normalization factor, denoted by $\hat{\sigma}_{d}(\ell)$, as:
    Choose a constant $\epsilon>0$, and let $\hat{\sigma}^{2}_{d,\epsilon}(\ell)=\max\{\hat{\sigma}^{2}_{d}(\ell),\epsilon\}$.
  6. 6.
  7. 7.
  8. 8.
    For $b=1,2,\cdots,B$, calculate the quantity
@@
Theorem 3.2 shows that the test $\phi_{n}$ has its size controlled asymptotically and is consistent. The proof for Theorem 3.2 is collected in Section B.2 of the online supplementary material. We also list all the technical assumptions, such as conditions that ensure the first-stage estimator converges at a sufficiently fast rate, in that section for the sake of exposition.
@@
Suppose Assumptions B.2, B.3, B.5 and B.1 in Section B.2 are satisfied. Let $\alpha\in(0,1/2)$ be the pre-chosen significance level.
@@
    $\limsup_{n\rightarrow\infty}\mathbb{P}(\phi_{n}=1|H_{1})=1.$ (3.10)
@@ 3.2 A semiparametric test with covariates dimension reduction
$Z\perp(Y_{0}(z),Y_{1}(z),D_{z};z\in\mathcal{Z})|X=x$ for all $x\in\mathcal{X}$.
@@
When Assumptions 3.1, 2.2 and 3.2 hold, the testable implications can be written as follows. For all $x\in\mathcal{X}$, $p_{1},p_{2}\in\mathcal{P}$ and $p_{1}\geq p_{2}$, and all $y,y^{\prime}\in\mathcal{Y}$
@@
If $X$ is discrete and $\mathcal{X}$ only contains a relatively small number of values, or $X$ contains a small number of continuous variables, we can also follow the same procedure as in Section 3.1 but add cubes for $X$. Under the null hypothesis, we should expect $T\equiv\sum_{d=0,1}\sum_{x\in\mathcal{X}}\sum_{\ell\in\mathcal{L}}\max\{\nu_{d}(\ell,x),0\}^{2}\Omega(\ell,x)=0$, where $\Omega(\ell,x)$ is a positive weighting function chosen by researchers, and
@@
$\nu_{0}(\ell,x)\equiv\mathbb{E}[(D-1)1\{y\leq Y\leq y+r_{y}\}1\{x\leq X\leq x+r_{x}\}1\{p_{2}\leq P\leq p_{2}+r_{p}\}]\\ \times\mathbb{E}[1\{x\leq X\leq x+r_{x}\}1\{p_{1}\leq P\leq p_{1}+r_{p}\}]-\mathbb{E}[1\{x\leq X\leq x+r_{x}\}1\{p_{2}\leq P\leq p_{2}+r_{p}\}]\\ \times{E}[(D-1)1\{y\leq Y\leq y+r_{y}\}1\{x\leq X\leq x+r_{x}\}1\{p_{1}\leq P\leq p_{1}+r_{p}\}],$
and $r_{x}$ is similarly defined as $r_{p}$ and $r_{y}$. The implementation follows analogously from 3.1.
@@
The advantage of using (3.13) and (3.14) is that both inequalities are only conditional on the scalar-valued propensity score. The effect of covariates has been filtered out by constructing a new outcome variable $\tilde{Y}$. Assumption 3.3 is a common assumption made in the literature for estimating the MTE, see for instance Carneiro and Lee (2009); Carneiro, Heckman, and Vytlacil (2010); Kowalski (2016). Nevertheless, we do acknowledge it is subject to the potential risk of model mis-specification. Under the null hypothesis of the model being correctly specified, parameters $\beta_{0}$ and $\beta_{1}$ can be estimated by partial linear regression of $Y$ on $X$ and propensity score $P$ separately for the sample of $D=1$ and $D=0$:
$\mathbb{E}[Y|X=x,P=p,D=d]=x^{\prime}\beta_{d}+K_{d}(p),\quad d\in\{0,1\},$
where $K_{d}(p)=\mathbb{E}[\alpha_{d}+U_{d}|X=x,D=d,P=p]$ only depends on $p$ under Assumption 3.3-(ii). The following algorithm summarizes the steps for implementation.
@@
[…]
@@
    Once $\hat{\beta}_{1}$ and $\hat{\beta}_{0}$ are obtained, one can construct estimates for ${\widetilde{Y}}_{i}=Y_{i}-X_{i}^{\prime}(D_{i}\hat{\beta}_{1}+(1-D_{i})\hat{\beta}_{0})$